# Explicit unit distance graphs with exponential chromatic number and arbitrary girth

### v1

@@ 3 Black-box arguments
In this section, we extract the main ideas used in the previous one and apply them together with state-of-the-art lower bounds on the chromatic number without any girth restrictions to prove our results.
The first one is that by doubling the number of dimensions one can embed an arbitrary blow-up of an orthogonality graph. Recall that for a graph $G$ and a positive integer $m$, the blow-up graph $G^{m}$ is the graph obtained from $G$ by replacing each vertex with an independent set of size $m$ and each edge with a $K_{m,m}$ subgraph between the two corresponding independent sets of size $m$.
@@
For any $m\in\mathbb{N}$ and orthogonality graph $G$ in $\mathbb{R}^{d}$, $G^{m}$ is an orthogonality graph in $\mathbb{R}^{2d}$.
@@
    Let $G$ be an orthogonality graph in $\mathbb{R}^{d}$ with representation $V\subseteq\mathbb{R}^{d}$. For each $\textbf{v}\in V$, let $\textbf{v}^{\prime}=(\textbf{v},\textbf{0})\in(\mathbb{R}^{d})^{2}=\mathbb{R}^{2d}$, and $\textbf{v}^{\prime\prime}=(\textbf{0},\textbf{v})\in(\mathbb{R}^{d})^{2}=\mathbb{R}^{2d}$. Then, for each orthogonal pair $\textbf{u},\textbf{v}\in V$, we have that $\operatorname{span}_{\mathbb{R}}\{\textbf{u}^{\prime},\textbf{u}^{\prime\prime}\}$ and $\operatorname{span}_{\mathbb{R}}\{\textbf{v}^{\prime},\textbf{v}^{\prime\prime}\}$ are orthogonal planes in $\mathbb{R}^{2d}$. For each $\textbf{v}\in V$, choose distinct points $\textbf{v}_{1},\ldots,\textbf{v}_{m}$ in $\operatorname{span}_{\mathbb{R}}\{\textbf{v}^{\prime},\textbf{v}^{\prime\prime}\}$ so that the points are also distinct between different $\textbf{v}\in V$. Now observe that for every orthogonal pair $\textbf{u},\textbf{v}\in V$, since $\operatorname{span}_{\mathbb{R}}\{\textbf{u}^{\prime},\textbf{u}^{\prime\prime}\}$ and $\operatorname{span}_{\mathbb{R}}\{\textbf{v}^{\prime},\textbf{v}^{\prime\prime}\}$ are orthogonal planes, we have for every $1\leq i,j\leq m$ that $\textbf{u}_{i}$ is orthogonal to $\textbf{v}_{j}$. Therefore, $G^{m}$ is an orthogonality graph in $\mathbb{R}^{2d}$. ∎
@@
Our argument is a modification of one of Nešetřil and Rödl [29] that they used to construct graphs with large girth and chromatic number, which itself is a modification of one of Tutte [8, 9] who constructed triangle-free graphs with large chromatic number. As part of our argument, we will need the existence of (explicit) hypergraphs with high girth and chromatic number. A cycle of length $\ell\geq 2$ in a hypergraph $H$ is a tuple $(F_{1},\ldots,F_{\ell})$ of distinct hyperedges such that there exist distinct elements $v_{1},\ldots,v_{\ell}\in V(H)$ with $v_{i}\in F_{i}\cap F_{i+1}$ for $11\leq i\leq\ell-$ 1, and $v_{\ell}\in F_{\ell}\cap F_{1}$. The girth of a hypergraph is equal to the length of a shortest cycle. The chromatic number of a hypergraph is defined as the minimum number of independent sets (sets not containing any edge fully) needed to partition its vertex set. The existence of hypergraphs of arbitrary uniformity with high girth and chromatic number follows from an immediate strengthening of the classical probabilistic argument of Erdős [12] and is due to Erdős and Hajnal [13]. Multiple explicit constructions have been found since see e.g. [1, 10, 14, 22, 26, 31, 36, 37]. Given a vertex $u$ of a graph $G$, we let $N_{G}(u)$ denote the neighbourhood of $u$ in $G$.
@@
For every graph $G$ and positive integer $g$ there exists a graph $G^{\prime}$ with girth at least $g$ and a surjective homomorpishm $h$ from $G^{\prime}$ to $G$ such that for any proper $k$-colouring $c^{\prime}$ of $G^{\prime}$, there exists a proper $k$-colouring $c$ of $G$ such that $c(v)\in c^{\prime}(h^{-1}(v))$ for all $v\in V(G)$. In particular, $\chi(G^{\prime})=\chi(G)$.
@@
    If $|V(G)|\leq 2$ the theorem is immediate since we may just take $G^{\prime}=G$ and let $h$ to be the identity. So, with this observation serving as a basis, we shall argue inductively. Let $u$ be a vertex of $G$, then we may assume that the theorem holds for $G-u$. Now, let $G_{u}$ be a graph with girth at least $g$ and $h_{u}$ a surjective homomorpishm from $G_{u}$ to $G-u$ such that for any proper $k$-colouring $c_{u}$ of $G_{u}$, there exists a proper $k$-colouring $c$ of $G-u$ such that $c(v)\in c_{u}(h_{u}^{-1}(v))$ for all $v\in V(G-u)$.
@@
    We construct our desired $G^{\prime}$ by taking a vertex disjoint union of all $G_{F}$ together with a disjoint independent set $U$ whose vertices we identify with $V(H)$. In addition, for each $F\in E(H)$, we add an arbitrary matching between the vertices of $U$ corresponding to those of $F$ and $h_{F}^{-1}(N_{G_{F}}(u))\subseteq V(G_{F})$. We define our desired $h:V(G^{\prime})\to V(G)$ by letting $h(v)=h_{F}(v)$ for all $v\in V(G_{F})$ and $h(v)=u$ for all $v\in U$. Since $h_{F}$ is a surjective homomorphism to $G-u$, so is $h$ to $G$.
@@
    Consider a cycle $C$ in $G^{\prime}$. If $C$ is contained in some $G_{F}$, then clearly $C$ has length at least $g$. Otherwise, if $C$ is not contained in some $G_{F}$, then $C$ must contain some vertices $u_{1},\ldots,u_{r}$ of $U$ which we label according to the order in which they appear on $C$ (choosing $u_{1}$ and direction arbitrarily). Observe that for each $1\leq r\leq i$, the internal vertices of the subpath of $C$ between $u_{i}$ and $u_{i+1}$ (taking $u_{r+1}=u_{1}$) are contained in a unique $G_{F_{i}}$, where $F_{i}\in E(H)$, with $u_{i},u_{i+1}\in F_{i}$. This follows since $u_{i}$ and $u_{i+1}$ can not be adjacent on the cycle (since they belong to $U$ which is independent) and since there are no edges between distinct $G_{F}$. Furthermore, for each $i$ we have that $F_{i}\not=F_{i+1}$ (again taking $F_{r+1}=F_{1}$), since the edge from $u_{i+1}$ to $G_{F_{i}}$ is unique and can be traversed only once by $C$. Since $H$ has girth at least $g$, it follows that $r\geq g$. Therefore, in either case, $C$ has length at least $g$. So, $G^{\prime}$ has girth at least $g$.
@@
    Consider some proper $k$-colouring $c^{\prime}$ of $G^{\prime}$. Restriction of $c^{\prime}$ to $U$ gives a $k$-coloring of $H$, which has chromatic number at least $k+1$, so some edge of $H$ must be monochromatic in this coloring. Namely, there is some colour $a$ and some $F\in E(H)$ such that $c^{\prime}(F)=\{a\}$. Therefore, since $c^{\prime}$ was proper and by our construction, there is a matching between $F\subseteq U$ and $h_{F}^{-1}(N_{G_{F}}(u))$ we know $c^{\prime}(h_{F}^{-1}(N_{G_{F}}(u)))\subseteq\{1,\ldots,k\}\backslash\{a\}$. So, by applying the inductive assumption to the restriction of $c^{\prime}$ to $G_{F}$ we obtain a proper $k$-colouring $c$ of $G-u$ such that $c(v)\in c^{\prime}(h^{-1}(v))$ for all $v\in V(G-u)$. Since $a\notin c^{\prime}(h_{F}^{-1}(N_{G_{F}}(u)))$ we can extend $c$ by letting $c(u)=a$ to obtain a desired proper $k$-coloring of $G$.
@@
    This in particular implies that if there exists a proper $k$-colouring of $G^{\prime}$ there also exists a proper $k$-coloring of $G$ so $\chi(G^{\prime})\geq\chi(G)$. On the other hand, since there is a homomorphism from $G^{\prime}$ to $G$, we have that $\chi(G^{\prime})\leq\chi(G)$, so $\chi(G^{\prime})=\chi(G)$. ∎
@@
Remark. In private communication Alon suggested an alternative way to prove Theorem 6 explicitly. The basic idea is to start with a high-girth Ramanujan graph (such as an LPS-graph [27]) on $|G|m$ vertices. One can then assign to each vertex of $G$ a disjoint set of $m$ of its vertices and remove any edges within parts or between parts corresponding to non-adjacent vertices of $G$. One can then show using a similar argument as we did with random subsampling that the chromatic number of this graph will remain the same as that of $G$, provided $m$ is chosen to be sufficiently large.
The final ingredient is the state-of-the-art construction of orthogonality graphs with large chromatic number due to Raigorodskii [42]. We note that the construction is similar in spirit to the one we presented in the previous section with the key additional idea being to use vectors with coordinates in $\{-1,0,1\}$ instead of just $\{-1,1\}$.
Theorem 8 (Raigorodskii [42]).
There exist orthogonality graphs in $\mathbb{R}^{d}$ with chromatic number at least $(1.2255+o(1))^{d}$.
@@
By combining Theorem 8 with Proposition 5 and Theorem 7, we obtain Theorem 3 (since $\sqrt{1.2255}=1.107...$). As discussed at the end of Section 2, Theorem 3 then implies Theorem 1 and 2.
@@ 4 Concluding remarks
In this paper, we proved that there exist, unit distance and orthogonality graphs in $\mathbb{R}^{d}$ and diameter graphs in $\mathbb{R}^{d^{2}}$ with chromatic number exponential in $d$ and with arbitrarily large girth.
@@
We note also that our Proposition 5 allows us to embed (faithfully) in $\mathbb{R}^{2d}$ a blow-up of any (faithful) orthogonality graph in $\mathbb{R}^{d}$. It could be interesting to determine how much we need to increase the dimension, if at all, for this to hold. One does need to increase the number of dimensions when $d\leq 3$ by the results on the classical Erdős Unit Distance Problem [15] but for $d\geq 4$ it might be possible to embed the blow-up in the same space. If so this would slightly improve the bases of exponentials in our results.
@@ Acknowledgements
We thank Noga Alon for sharing his alternative explicit proof of Theorem 6 and thank Jim Geelen for helpful discussions on topics related to Theorem 6.

### v3

@@ 3 Black-box arguments
In this section, we extract the main ideas used in the previous one. We state them in a rather general form to make them more readily available for use in the future. Finally, we combine them together with the state-of-the-art lower bounds on the chromatic number without any girth restrictions to prove our main results.
The first ingredient is the state-of-the-art construction of orthogonality graphs with large chromatic number due to Raigorodskii [46]. We note that the construction is similar in spirit to the one we presented in the previous section with the key additional idea being to use vectors with coordinates in $\{-1,0,1\}$ instead of just $\{-1,1\}$.
Theorem 5 (Raigorodskii [46]).
There exist orthogonality graphs in $\mathbb{S}^{d-1}$ with chromatic number at least $(2/\sqrt{3}+o(1))^{d}\geq(1.1547+o(1))^{d}$.
The second ingredient is the fact that by doubling the number of dimensions one can embed an arbitrary blow-up of an orthogonality graph. Recall that for a graph $G$ and a positive integer $m$, the blow-up graph $G^{m}$ is the graph obtained from $G$ by replacing each vertex with an independent set of size $m$ and each edge with a $K_{m,m}$ subgraph between the two corresponding independent sets of size $m$.
@@
For any $m\in\mathbb{N}$ and orthogonality graph $G$ in $\mathbb{S}^{d-1}$, $G^{m}$ is an orthogonality graph in $\mathbb{S}^{2d-1}$.
@@
    Let $G$ be an orthogonality graph in $\mathbb{S}^{d-1}\subseteq\mathbb{R}^{d}$ with representation $V\subseteq\mathbb{S}^{d-1}$. For each $\textbf{v}\in V$, let $\textbf{v}^{\prime}=(\textbf{v},\textbf{0})\in\mathbb{S}^{d-1}\times\mathbb{R}^{d}\subseteq\mathbb{R}^{2d}$, and $\textbf{v}^{\prime\prime}=(\textbf{0},\textbf{v})\in\mathbb{R}^{d}\times\mathbb{S}^{d-1}\subseteq\mathbb{R}^{2d}$. Then, for each orthogonal pair $\textbf{u},\textbf{v}\in V$, we have that $\operatorname{span}_{\mathbb{R}}\{\textbf{u}^{\prime},\textbf{u}^{\prime\prime}\}$ and $\operatorname{span}_{\mathbb{R}}\{\textbf{v}^{\prime},\textbf{v}^{\prime\prime}\}$ are orthogonal planes in $\mathbb{R}^{2d}$. For each $\textbf{v}\in V$, choose distinct points $\textbf{v}_{1},\ldots,\textbf{v}_{m}$ of unit length in $\operatorname{span}_{\mathbb{R}}\{\textbf{v}^{\prime},\textbf{v}^{\prime\prime}\}$ so that the points are also distinct between different $\textbf{v}\in V$. Now observe that for every orthogonal pair $\textbf{u},\textbf{v}\in V$, since $\operatorname{span}_{\mathbb{R}}\{\textbf{u}^{\prime},\textbf{u}^{\prime\prime}\}$ and $\operatorname{span}_{\mathbb{R}}\{\textbf{v}^{\prime},\textbf{v}^{\prime\prime}\}$ are orthogonal planes, we have for every $1\leq i,j\leq m$ that $\textbf{u}_{i}$ is orthogonal to $\textbf{v}_{j}$. Therefore, $G^{m}$ is an orthogonality graph in $\mathbb{S}^{2d-1}$. ∎
@@
Our argument is a modification of one of Nešetřil and Rödl [30] that they used to construct graphs with large girth and chromatic number, which itself is a modification of one of Tutte [9, 10] (using the pseudonym Blanche Descartes) who constructed triangle-free graphs with large chromatic number. As part of our argument, we will need the existence of (explicit) hypergraphs with high girth and chromatic number. A cycle of length $\ell\geq 2$ in a hypergraph $H$ is a tuple $(F_{1},\ldots,F_{\ell})$ of distinct hyperedges such that there exist distinct elements $v_{1},\ldots,v_{\ell}\in V(H)$ with $v_{i}\in F_{i}\cap F_{i+1}$ for $11\leq i\leq\ell-$ 1, and $v_{\ell}\in F_{\ell}\cap F_{1}$. The girth of a hypergraph is equal to the length of a shortest cycle. The chromatic number of a hypergraph is defined as the minimum number of independent sets (sets not containing any edge fully) needed to partition its vertex set. The existence of hypergraphs of arbitrary uniformity with high girth and chromatic number follows from an immediate strengthening of the classical probabilistic argument of Erdős [13] and is due to Erdős and Hajnal [14]. Multiple explicit constructions have been found since see e.g. [1, 11, 15, 22, 27, 32, 38, 37]. Given a vertex $u$ of a graph $G$, we let $N_{G}(u)$ denote the neighbourhood of $u$ in $G$. A homomorphism from a graph $G^{\prime}$ to a graph $G$ is a function from $V(G^{\prime})\to V(G)$ such that if $vu\in E(G^{\prime})$ then $h(v)h(u)\in E(G)$. Note that there exists a homomorphism from $G^{\prime}$ to $G$ if and only if $G^{\prime}$ is a subgraph of a blow-up of $G$.
@@
For every graph $G$ and pair of positive integers $g,k$ there exists a graph $G^{\prime}$ with girth at least $g$ and a surjective homomorpishm $h$ from $G^{\prime}$ to $G$ such that for any proper $k$-colouring $c^{\prime}$ of $G^{\prime}$, there exists a proper $k$-colouring $c$ of $G$ such that $c(v)\in c^{\prime}(h^{-1}(v))$ for all $v\in V(G)$. In particular, $\chi(G^{\prime})=\chi(G)$.
@@
    If $|V(G)|\leq 2$ the theorem is immediate since we may just take $G^{\prime}=G$ and let $h$ be the identity. So, with this observation serving as a basis, we shall argue by induction on the number of vertices (while keeping $g$ and $k$ fixed). Let $u$ be a vertex of $G$, then we may assume that the theorem holds for $G-u$. Now, let $G_{u}$ be a graph with girth at least $g$ and $h_{u}$ a surjective homomorphism from $G_{u}$ to $G-u$ such that for any proper $k$-colouring $c_{u}$ of $G_{u}$, there exists a proper $k$-colouring $c$ of $G-u$ such that $c(v)\in c_{u}(h_{u}^{-1}(v))$ for all $v\in V(G-u)$.
@@
    We construct our desired $G^{\prime}$ by taking a vertex disjoint union of all $G_{F}$ together with a disjoint independent set $U$ whose vertices we identify with $V(H)$. In addition, for each $F\in E(H)$, we add an arbitrary matching between the vertices of $U$ corresponding to those of $F$ and $h_{F}^{-1}(N_{G}(u))\subseteq V(G_{F})$. We define our desired $h:V(G^{\prime})\to V(G)$ by letting $h(v)=h_{F}(v)$ for all $v\in V(G_{F})$ and $h(v)=u$ for all $v\in U$. Since $h_{F}$ is a surjective homomorphism to $G-u$, so is $h$ to $G$. For an illustration of this construction, see Figure 1.
    Figure 1: An illustration of the construction for $G^{\prime}$ when $G$ is the Moser spindle. The labels of the vertices of $G_{u}$ gives the homomorphism from $G_{u}$ to $G-u$. For simplicity, we only show how one copy $G_{F}$ of $G_{u}$ attaches to vertices $F$ of $U$ in $G^{\prime}$. In reality, we place one such copy for every edge of $H$.
@@
Consider a cycle $C$ in $G^{\prime}$. If $C$ is contained in some $G_{F}$, then clearly $C$ has length at least $g$. Otherwise, if $C$ is not contained in some $G_{F}$, then $C$ must contain some vertices $u_{1},\ldots,u_{r}$ of $U$ which we label according to the order in which they appear on $C$ (choosing $u_{1}$ and direction arbitrarily). Observe that for each $1\leq i\leq r$, the internal vertices of the subpath of $C$ between $u_{i}$ and $u_{i+1}$ (taking $u_{r+1}=u_{1}$) are contained in a unique $G_{F_{i}}$, where $F_{i}\in E(H)$, with $u_{i},u_{i+1}\in F_{i}$. This follows since $u_{i}$ and $u_{i+1}$ can not be adjacent on the cycle (since they belong to $U$ which is independent) and since there are no edges between distinct $G_{F}$. Furthermore, for each $i$ we have that $F_{i}\not=F_{i+1}$ (again taking $F_{r+1}=F_{1}$), since the edge from $u_{i+1}$ to $G_{F_{i}}$ is unique and can be traversed only once by $C$. Therefore, $H$ has a cycle consisting of (not necessarily all of) the edges $F_{1},\ldots,F_{r}$. Since $H$ has girth at least $g$, it follows that $r\geq g$. Therefore, in either case, $C$ has length at least $g$. So, $G^{\prime}$ has girth at least $g$.
@@
Consider some proper $k$-colouring $c^{\prime}$ of $G^{\prime}$. Restriction of $c^{\prime}$ to $U$ gives a $k$-coloring of $H$, which has chromatic number at least $k+1$, so some edge of $H$ must be monochromatic in this coloring. Namely, there is some colour $a$ and some $F\in E(H)$ such that $c^{\prime}(F)=\{a\}$. Therefore, since $c^{\prime}$ was proper, and by our construction there is a matching between $F\subseteq U$ and $h_{F}^{-1}(N_{G}(u))$, we conclude $c^{\prime}(h_{F}^{-1}(N_{G}(u)))\subseteq\{1,\ldots,k\}\backslash\{a\}$. So, by applying the inductive assumption to the restriction of $c^{\prime}$ to $G_{F}$ we obtain a proper $k$-colouring $c$ of $G-u$ such that $c(v)\in c^{\prime}(h^{-1}(v))$ for all $v\in V(G-u)$. Since $a\notin c^{\prime}(h_{F}^{-1}(N_{G}(u)))$ we can extend $c$ by letting $c(u)=a$ to obtain a desired proper $k$-coloring of $G$.
@@
This, in particular, implies that if there exists a proper $k$-colouring of $G^{\prime}$, there also exists a proper $k$-coloring of $G$, so $\chi(G^{\prime})\geq\chi(G)$. On the other hand, since there is a homomorphism from $G^{\prime}$ to $G$, we have that $\chi(G^{\prime})\leq\chi(G)$, so $\chi(G^{\prime})=\chi(G)$. ∎
@@
Remark. In private communication, Alon suggested an alternative way to prove Theorem 7 explicitly. The basic idea is to start with a high-girth Ramanujan graph (such as an LPS-graph [28]) on $|G|m$ vertices. One can then assign to each vertex of $G$ a disjoint set of $m$ of its vertices and remove any edges within parts or between parts corresponding to non-adjacent vertices of $G$. One can then show, using a similar argument as we did with random subsampling, that the chromatic number of this graph will remain the same as that of $G$, provided $m$ is chosen to be sufficiently large.
@@
By combining Theorem 5 with Proposition 6 and Theorem 8, we obtain Theorem 3 (since $\sqrt{1.1547}=1.074569...$). As discussed at the end of Section 2, Theorem 3 then implies Theorem 1 and 2 (where $1.074569^{\sqrt{2}}=1.107$).
@@ 4 Concluding remarks
In this paper, we proved that there exist unit distance and orthogonality graphs in $\mathbb{R}^{d}$ and diameter graphs in $\mathbb{R}^{d^{2}}$ with chromatic number exponential in $d$ and with arbitrarily large girth.
@@
We note also that our Proposition 6 allows us to embed (faithfully) in $\mathbb{S}^{2d-1}$ a blow-up of any (faithful) orthogonality graph in $\mathbb{S}^{d-1}$. This is in general tight since one can embed a complete graph $K_{d}$ as an orthogonality graph in $\mathbb{S}^{d-1}$ but to embed a complete $d$-partite graph with three vertices in every part onto a sphere one needs at least $2d$ dimensions. Indeed, each triple of vertices making a part defines a plane which should be orthogonal to all other such planes. We thank an anonymous referee for pointing this out.
@@ Acknowledgements
We thank Noga Alon for sharing his alternative explicit proof of Theorem 7 and thank Jim Geelen for helpful discussions on topics related to Theorem 7. We also thank Andrei M. Raigorodskii for pointing out that the orthogonality graphs he constructs in [46] come with a slightly worse base constant than we initially stated. We are grateful to the anonymous referees for their careful reading of the paper and many useful suggestions.